OpenStudy (viixxen):

"Is 16, 34, 30 a Pythagorean Triple? Explain your answer." I think it is, but I'm not completely sure. Thank you! (:

OpenStudy (campbell_st):

well you can check is \[34^2 = 30^2 + 16^2\] that's the rule it needs to follow to be a pythagorean triple

OpenStudy (adi3):

@campbell_st so tht means it is right

OpenStudy (campbell_st):

well providing its true... I don't have a calculator 16^2 = 256 30^2 = 900 not sure what 34^2... but if the sum of the 2 smaller squares equals the larger square, you have a triple.

OpenStudy (hhopke):

Yes, because 16^2 + 30^2 = 34^2 (256+900=1156)